A population has mean 73.5 and standard deviation 2.5. Find the mean and standard deviation of X ̅ for samples of size 30.


1
Expert's answer
2022-05-17T08:19:45-0400
"\mu_{\bar{X}}=\mu=73.5"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{2.5}{\sqrt{30}}\approx0.4564"